# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

### v3

@@ 4.1 Results and Discussions
Models trained with $\mathsf{Inheritune}$ outperform much larger models trained from scratch.
@@
We present our main results in Figure 5. The 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium, respectively, achieve comparable or lower validation losses than both their full-sized counterparts and same-sized models trained from scratch, when trained for the same number of steps (100K). Our GPT-2 xlarge and GPT-2 large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations. From a convergence perspective, prior work has linked overparameterization to faster convergence Bengio et al. (2005); Vaswani et al. (2018). Interestingly, we find that smaller models derived using $\mathsf{Inheritune}$ converge just as fast as their larger counterparts. A detailed summary of these results is provided in the appendix Table 3.
@@
We conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023); Biderman et al. (2023) using the high-quality Fineweb_edu dataset. We trained a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following $\mathsf{Inheritune}$. For comparison, we also trained 16-layer baseline models from scratch. All models were trained for 100K steps. We observe thematically consistent results: as shown in Figure 6, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation. Downstream results are averaged over five tasks (see table within Figure 6); the complete breakdown is provided in Table 5 placed in appendix. Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
@@
Models trained with Inheritune outperform all zero-shot model initialization baselines.
@@
In Table 1, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 14.
Distillation vs Inheritune.
In Figure 7, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K training steps. Additional distillation experiments can be found in the supplementary materials.
4.2 Ablations
Layers Initialization Steps Val loss ($\downarrow$)
16 attn (w/ layernorm) 100K 2.84
16 mlp (w/ layernorm) 100K 2.85
16 attn+mlp (w/o layernorm) 100K 2.80
16 Ours (attn+mlp w/ layernorm) 100K 2.81
Table 2: Impact of initializing various sub-modules within a transformer block. We compare validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 15.
We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence. For these ablations, we fixed the model to a 16-layer GPT-2 medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 medium reference model. We initialize the transformer blocks with 1) attention ((key, query, value, and projection) and the layernorm [^2] In GPT-2 models layernorm blocks are parameterized. weights (attn w/ layernorm), 2) attention and mlp weights without the layer-norm (attn+mlp w/o layernorm), and 3) mlp weights with the layer norm (mlp w/ layernorm). We emphasize that $\mathsf{Inheritune}$ performs initialization by inheriting attention and mlp weights with the layer norm (attn+mlp w/ layernorm).
As shown in Table 2, models trained with attention and mlp weights demonstrated the best performance, regardless of the layer norm initialization. A detailed validation loss vs training steps plot is presented in supplementary Figure 15. We conclude that initializing both attention and MLP weights provides a clear advantage. Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
5 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied is a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). Recently He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023) employs a stage-wise training strategy that uses weights from initial layers to initialize later layers has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
6 Conclusion
In this paper, we identified a structural flaw in the attention mechanism of deep decoder-style LLMs, where many deeper layers tend to lose rank and converge into near rank-1 matrices. To address this, we propose $\mathsf{Inheritune}$, to train smaller models that inherits early blocks from a larger model and expands the architecture gradually, matching the performance of the reference model. We validated $\mathsf{Inheritune}$ on GPT-2 models of varying sizes, achieving efficient smaller models without performance loss on the OpenWebText and FineWeb_edu datasets.
Supplementary Materials
Contents
  • A: Extended Discussion on Attention Collapse
  • B: Baselines
  • C: Supplementary Experiments
  • D: Extended Discussion
  • E: Additional Experiments on Low Data Regime
  • F: Implementation Details
Appendix A Extended Discussion on Attention Collapse
A.1 Attention Mass Analysis
Dominant Single-Column Structure in Attention Matrices.
In Figure 8, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models. For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer. We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$. We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion. As shown in Figure 9, we performed a mass analysis on contemporary billion-parameter LLaMA models (Geng & Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers. This provides concrete evidence that the phenomenon persists across a broad range of architectures at the billion-parameter scale.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 xLarge (1.5B) L=48, H=1600
(a) OpenLLaMA-3B
(b) OpenLLaMA-7B
(c) OpenLLaMA-13B
Figure 9: The overall mass of attention matrices in billion-scale LLMs, pre-trained on trillions of tokens, tends to concentrate in fewer columns. This phenomenon becomes increasingly pronounced as the model size grows. We computed attention matrices using 100 tokens from a random subset of RedPajama with 1B tokens. Next, we performed 100 runs and plotted the mean and standard deviation of the mass as a function of layers for our mass analysis, respectively. We followed the same procedure as discussed in Section 2. Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng & Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized. Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column. This observation aligns closely with our analysis in Figure 1.
A.2 Additional Analysis on Attention Rank Collapse.
To investigate the robustness of our analysis, we performed a rank analysis on GPT-2 models trained on OpenWebText (as shown in Figure 1) using a validation set from FineWeb-Edu—a new dataset on which the models were not trained (see Figure 10). It is clearly observed that attention consistently loses rank, often collapsing to single-column structures across all models, thereby reinforcing the robustness of our findings.
Next, we conducted a rank analysis on a contemporary LLaMA-3 model with 8B parameters. We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019). As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red. This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
A.3 Attention Pattern Visualization
To provide further evidence of lazy layers, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 12 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models computes causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over 3 runs.
@@ A.3 Attention Pattern Visualization
In the 24-layer model trained from scratch (top row of Fig. 12), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes model more attentive and addresses attention degeneration, potentially leading to more efficient models in compact size (also refer Figure 17 and Figure 18). Additionally, in Figure 13, we present attention patterns from a 48-layer GPT-2 xLarge model, comparing attention maps from a potent (non-lazy) and a lazy layer.
(a) Rank analysis of GPT-2 Medium
(b) Rank analysis of GPT-2 Large
(c) Rank analysis of GPT-2 XLarge
(d) Matrix mass analysis of GPT-2 Medium
(e) Matrix mass analysis of GPT-2 Large
(f) Matrix mass analysis of GPT-2 XLarge
Figure 10: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations. All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the Fineweb_edu validation set, following the same procedure described in Figure 1. This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
@@
24 layer GPT2 medium model trained from scratch.
@@
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
@@
Figure 12: $\mathsf{Inheritune}$ preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected string, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s deeper layers.

### v4

@@ 4.1 Results and Discussions
Models trained with Inheritune outperform both larger and same sized models trained from scratch.
@@
We present our main results in Figure 5. The 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 XLarge, 36-layer GPT-2 Large, and 24-layer GPT-2 Medium, respectively, achieve comparable or lower validation losses than both their full-sized counterparts and same-sized models trained from scratch, when trained for the same number of steps (100K). Our GPT-2 XLarge and GPT-2 Large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 Medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations. Furthermore, as shown in Figure 6, models trained with $\mathsf{Inheritune}$ reach the same validation loss as same-sized models trained from scratch in approximately half the number of training steps. Moreover, for the GPT-2 Medium and Large variants, $\mathsf{Inheritune}$ achieves a strictly lower loss floor that same-sized models fail to reach even when trained for twice as many steps. A tabular summary of these results is provided in Appendix Table 5.
@@
We conducted additional training experiments, using a high-quality training data namely Fineweb. We trained a custom 32-layer GPT-2 Large (668M) and a 24-layer GPT-2 Medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 Large and the other from GPT-2 Medium, using their respective reference models following $\mathsf{Inheritune}$. For comparison, we also trained 16-layer baseline models from scratch. All models were trained for 100K steps, and training loss was used to evaluate pre-training performance. We observe thematically consistent results: as shown in Figure 7, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation. Downstream results are provided in Table 1. Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
@@
Models trained with Inheritune outperform all warm started baselines.
@@
In Table 2, we compare GPT-2 XLarge, GPT-2 Large, and GPT-2 Medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid stacking, and half-width baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy, with a single exception in the GPT-2 XLarge case where it matches one baseline. For a detailed view of the training curves across all methods, refer to the training curves in supplementary Figure 18.
4.2 Inheritune Mitigates Attention Collapse
We attribute the success of $\mathsf{Inheritune}$, to its ability to mitigate attention collapse, thereby leading to fewer lazy layers after training. In Figures 8 we juxtapose the attention rank patterns of the vanilla and $\mathsf{Inheritune}$-trained GPT-2 Medium. Notably, none of the GPT-2 Medium variants exhibit lazy layers. A similar analysis is conducted with GPT-2 XLarge (refer supplementary Figure 13).
@@ 4.2 Inheritune Mitigates Attention Collapse
The corresponding attention patterns for GPT-2 Medium, shown in Figure 9, further corroborate our observation. The attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models compute causal attention. We computed these attention matrices using randomly selected strings from the validation set of OpenWebText and took 40 tokens averaged over 3 runs. In the 24-layer model trained from scratch (top row of Figure  9), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit dense structured attention patterns. In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus (attention). This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes the model more attentive and addresses attention collapse, potentially leading to more efficient models in compact size.
Figure 8: Rank collapse in deeper layers and its mitigation through Inheritune. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Figure 1. Analysis of a 24-layer GPT-2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs). Our 16-layer GPT-2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
@@
Vanilla 24 layer GPT-2 Medium
@@
GPT-2 Medium 16 layer variant (Ours)
@@
Figure 9: Inheritune preserves effective attention patterns. Comparison of attention patterns across layers (L) and heads (H) in two GPT-2 medium models: (top) vanilla 24-layer model trained from scratch, (bottom) 16-layer variant trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the vanilla model’s deeper layers.
4.3 Ablations and Limitations
We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence. For these ablations, we fixed the model to a 16-layer GPT-2 Medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 Medium reference model. We initialize the transformer blocks with 1) Attention block (key, query, value, and projection weights) along with the layernorm [^2] In GPT-2 models layernorm blocks are parameterized., 2) Attention and MLP (FFN) weights without the layer-norm weights, and 3) MLP block weights along with the layernorm. We note that $\mathsf{Inheritune}$ performs initialization by inheriting the entire layer weight i.e. attention, MLP along with the layernorm weights.
As shown in Table 3, models initialized with both Attention and MLP weights achieve the best performance, irrespective of the LayerNorm initialization. A detailed validation loss versus training steps plot is provided in the supplementary Figure 19. These results suggest that jointly initializing the Attention and MLP submodules offers a clear advantage over initializing either component alone. Interestingly, we also find that initializing only the Attention or only the MLP weights yields comparable improvements in both convergence speed and final validation loss.
Sensitivity Analysis for attention rank computation and lazy layers.
Next, we analyze the sensitivity of the approximate rank to the variance threshold $\tau$. Following the rank computation methodology described in Section 2, we perform rank analysis on the full GPT-2 Medium and GPT-2 Large models using a randomly sampled subset of the OpenWebText’s validation set, varying $\tau\in\{0.8,0.85,0.9,0.95\}$. As shown in Figure 10, the $\mathrm{MaxRank}(l)$ (y-axis) remains highly stable for the lazy layers across all values of $\tau$ for both the models.
Limitations of our work.
Our analysis primarily focuses on pre-layernorm (Pre-LN) architectures. By default, we initialize training at $l=L/2$, i.e. from the midpoint of the model where early layers have already developed strong representations to minimize the total number of training rounds. Figure 20 presents an ablation study showing training curves across three rounds performed following our proposed method. Notably, by the third round ($R=3$), models trained with $\mathsf{Inheritune}$ match the validation loss of their reference counterpart (Full model). Despite its effectiveness, $\mathsf{Inheritune}$ is computationally expensive, as it may require multiple rounds of training during the growth phase. Lastly, our current analysis focuses exclusively on attention submodules; extending this framework to feedforward (MLP) layers remains an important direction for future research.
Layers Initialization Steps Pre-train Val loss ($\downarrow$)
16 Attention 100K 2.84
16 MLP 100K 2.85
16 Attention and MLP* 100K 2.80
16 Ours 100K 2.81
Table 3: Impact of initializing various sub-modules within a transformer block. We compare validation loss of a 16-layer GPT-2 Medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 Medium reference model. Submodules without the * marker also include layernorm weights. All models are trained with the OpenWebText dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and Attention and MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. The training curves of corresponding models are presented in Figure 19. The lowest and our corresponding validation losses (lower is better) are highlighted in bold.
(a) $\tau$=0.80
(b) $\tau$=0.85
(c) $\tau$=0.90
(d) $\tau$=0.95
(e) $\tau$=0.80
(f) $\tau$=0.85
(g) $\tau$=0.90
(h) $\tau$=0.95
Figure 10: Lazy layers remain robust under variations of $\tau$. To study the sensitivity of attention rank to the variance–ratio threshold $\tau\in\{0.8,0.85,0.9,0.95\}$, we visualize $\mathrm{MaxRank}(l)$ as a function of layer index for (a–d) the GPT-2 Medium full model with 24 layers (top row) and (e–h) the GPT-2 Large full model with 36 layers (bottom row).
5 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied is a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024); Wu et al. (2024). Recently He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two different model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate collapse. In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023) employs a stage-wise training strategy that uses weights from initial layers to initialize later layers has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs. In some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman and Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
6 Conclusion
In this work, we identify a structural inefficiency in deep decoder-style LLMs, which we term attention collapse, where attention matrices in deeper layers often degenerate into near rank-one structures, rendering these layers ineffective. These ineffective layers referred to as lazy layers contribute little to the model’s representational power. To address this, we introduce $\mathsf{Inheritune}$, a multi-stage training recipe that initializes a smaller model using a few potent early layers from a larger pre-trained model and then progressively trains and expands it through multiple rounds. Our experiments demonstrate that models trained with $\mathsf{Inheritune}$ can match or even surpass the performance of their larger counterparts despite having significantly fewer layers. By mitigating attention collapse, our approach produces compact and highly performant models, offering a new path toward designing smaller, more attentive architectures from the ground up.
Acknowledgments
This research was supported by NSF Grant 2217069 and by UT Austin’s Center for Generative AI and the Machine Learning Lab. We thank Nived Rajaraman for helpful discussions and the anonymous reviewers for their valuable feedback.
Supplementary Materials
Contents
  • A: Frequently Asked Questions
  • B: Extended Discussion on Attention Collapse
  • C: Understanding Attention Degradation using Attention Mass Analysis
  • D: Baselines
  • E: Supplementary Experiments
  • F: Model Configurations and Training Details
  • G: Additional Experiments and Discussions
  • H: Theoretical Analysis
Appendix A Frequently Asked Questions
A.1 Is your method not just pruning?
$\mathsf{Inheritune}$ is a stage-wise efficient training recipe that addresses a structural issue in decoder-style transformer blocks—Attention collapse—which we consistently observed across multiple models.
Unlike pruning, Inheritune includes a growth phase where the model is expanded until it outperforms the reference model (refer Algorithm 1 and Figure 20). Pruning doesn’t always require re-training (Ma et al., 2023), whereas $\mathsf{Inheritune}$ may need multiple rounds of re-training. To the best of our knowledge, no pruning method has explicitly studied or resolved attention collapse in LLMs. Our method has closer proximity to efficient training recipes employing model initialization (J. Reddi et al., 2023; Du et al., 2024) or warm starting (Ash and Adams, 2020).
A.2 Is the comparison with baseline models trained from scratch unfair since Inheritune uses weights from pre-trained models?
In Baseline-II (refer Section 4) we have also compared our method with warm started baselines which also uses pre-trained model weights for model initialization for fair comparisons.
For Baseline-I, we include much larger reference models as well as same-sized models trained for twice as many steps. Remarkably, $\mathsf{Inheritune}$ still outperforms both. We believe these findings are novel and reveal a new axis for scaling.
A.3 The attention collapse analysis is not holistic?
TL;DR.
We analyzed the phenomenon of attention collapse across four datasets and four different model architectures, with model sizes ranging from millions to billions of parameters.
We evaluated attention patterns and analyzed the phenomenon of attention collapse using four datasets: OpenWebText, FineWeb, RedPajama, and C4. Our analysis used two complementary metrics namely, approximate rank and approximate mass to quantify the structure of attention matrices. We conducted attention collapse analysis on a range of models, including GPT-2 (Medium 355M, Large 770M, XLarge 1.5B), LLaMA-3 (3B and 8B models ), OLMo 1B, Cerebras-GPT 2.7B, Falcon 7B and LLaMA-1 (OpenLLaMA 3B, 7B, 13B) which features a notably different architecture. The evaluation was performed on 10K tokens (100 samples $\times$ 100 tokens each), and we also provide visualizations of the resulting attention patterns.
A.4 What is the connection between Attention collapse and Attention sinks?
The term Attention sink (Xiao et al., 2024) refers to a phenomenon where a specific token in a sequence receives disproportionately high attention scores compared to other tokens in the attention map.
In our analysis of Attention collapse, we also observed sink-like behavior for certain tokens across all attention maps (see Figure 9 and Figure 12). However, unlike typical attention sinks, we found that beyond the sink token, no meaningful attention structure remains: all other tokens receive nearly uniform attention scores. We further connect this behavior to the emergence of lazy layers. Therefore our analysis has unique insights compared to attention sinks.
Appendix B Extended Discussion on Attention Collapse
B.1 Attention Collapse in LLaMA-3 models
We conducted a rank analysis on a contemporary LLaMA-3 base model with 8B parameters. We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019). As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red. This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy. A different pattern of attention collapse compared to GPT-2 models can be attributed to the architectural differences between these models.
B.2 Attention Pattern Visualization of Potent and Lazy Layers
Following the analysis in Section 2 and Figure 1, we present the attention patterns of two representative layers from the GPT-2 XL model: a lazy layer (Layer 30) and a potent layer (Layer 8), as shown in Figure 12. The attention patterns of these layers exhibit distinctly different behaviors. In particular, the lazy layer demonstrates a clear collapse, where attention concentrates almost exclusively on the first token. Next, following the discussions in Section 4.2, we visualize the rank–layer relationship for the GPT-2 XLarge model, juxtaposing a vanilla model with a $\mathsf{Inheritune}$-trained model containing half as many layers. Although both models achieve similar validation loss, the $\mathsf{Inheritune}$-trained model exhibits significantly fewer lazy layers compared to the vanilla counterpart.